OpenStudy (anonymous):

1.2258 = 9.8(sin θ) I do not know how to solve trigonometric equations and would really appreciate any help!:) I need to solve for theta.

OpenStudy (anonymous):

\[\large \sf 1.2258=9.8sin(\theta)\]\[\large \sf \frac{1.2258}{9.8}=sin(\theta)\]\[\large \sf asin \huge ( \large \frac{1.2258}{9.8} \huge) \large=\theta\]

OpenStudy (anonymous):

@amorejon